# Uncertainty-Aware Deep Learning Framework for Remaining Useful Life Prediction in Turbofan Engines with Learned Aleatoric Uncertainty

### v1

@@ V-C Learned Uncertainty Quantification Analysis
V-D Performance Stratification by RUL Range
TABLE III: RMSE by RUL Range (FD001 Best Run)
@@ V-D Performance Stratification by RUL Range
Table III reveals performance stratification across degradation phases. As expected, prediction accuracy decreases in early operational life where degradation signatures are subtle and sensor readings remain near nominal values. However, as engines approach failure (RUL $\leq$ 30), our framework achieves exceptional accuracy (RMSE 5.14), precisely where accuracy matters most for preventing catastrophic failures while avoiding unnecessary premature maintenance.
@@
V-E Multi-Condition Dataset Performance
@@
V-F Ablation Study and Component Analysis
@@
V-G Computational Efficiency Analysis
@@ V-H Comparative Discussion
Computational Complexity vs. Deployment Feasibility: Transformer-based methods achieving 11-12 RMSE require 2M+ parameters and substantial computational resources. Domain-adaptation approaches like TMSCNN, while achieving excellent results, employ complex transfer learning pipelines. Our 487K-parameter architecture achieves ¡10ms inference on M1 hardware, making it practical for real-time onboard prognostics without ground station dependencies.
@@
V-I Practical Implications for Maintenance Scheduling

### v2

@@ V-C Learned Uncertainty Quantification Analysis
V-D False Positive and False Negative Analysis for Maintenance Decisions
For practical predictive maintenance deployment, understanding false positives (unnecessary maintenance warnings) and false negatives (missed critical maintenance needs) is as crucial as overall RMSE. We define maintenance decision thresholds where engines with predicted RUL $\leq$ 30 cycles require immediate maintenance intervention, while those exceeding this threshold are considered healthy for continued operation.
TABLE III: Maintenance Decision Analysis (FD001, Threshold: RUL $\leq$ 30)
Metric Value
True Positives (TP) 24
True Negatives (TN) 72
False Positives (FP) 3
False Negatives (FN) 1
Accuracy 96.0%
Precision 88.9%
Recall (Sensitivity) 96.0%
F1 Score 92.3%
False Positive Rate 4.0%
False Negative Rate 4.0%
True Positive Rate 96.0%
Specificity 96.0%
Table III presents maintenance decision classification metrics on FD001. The model achieves 96% accuracy with balanced error rates: only 3 false alarms (4.0% FPR) representing unnecessary maintenance interventions, and critically, just 1 missed detection (4.0% FNR) out of 25 engines requiring maintenance. This 96% recall rate for critical maintenance needs demonstrates excellent safety characteristics.
Figure 3: False positive and false negative analysis for maintenance decisions on FD001. Top-left: Confusion matrix at RUL $\leq$ 30 threshold showing excellent classification with only 1 false negative. Top-right: FPR and FNR across varying RUL thresholds, demonstrating optimal operating point around 30 cycles. Bottom-left: Precision-recall trade-off showing robust F1 scores (92%) across thresholds. Bottom-right: Error distribution revealing conservative model behavior with 65% under-predictions (safer) versus 35% over-predictions, ideal for safety-critical applications.
Figure 3 provides comprehensive FP/FN visualization. The confusion matrix demonstrates strong classification performance with minimal false negatives—critical for preventing catastrophic failures. Analysis across multiple thresholds (20-100 cycles) reveals that our chosen threshold of 30 cycles achieves optimal balance, with both FPR and FNR at 4%.
The error distribution analysis reveals favorable characteristics: 65% of predictions are under-predictions (predicting lower RUL than actual, mean error 15.5 cycles) versus 35% over-predictions (mean error 16.0 cycles). This conservative bias is ideal for safety-critical maintenance—the model errs toward earlier maintenance warnings rather than delayed interventions. Over-predictions represent the dangerous scenario where predicted RUL exceeds actual remaining life, potentially delaying necessary maintenance. Our framework limits these to 35 cases (35%) with maximum over-prediction of 50.3 cycles, acceptable given the learned uncertainty bounds provide additional safety margins.
Cost-benefit implications are clear: 3 false positives represent manageable unnecessary maintenance interventions (wasted resources, downtime), while 1 false negative poses safety risk requiring secondary inspection protocols. The 96% detection rate for critical maintenance needs, combined with calibrated uncertainty estimates, enables risk-aware scheduling where high-uncertainty predictions near thresholds trigger additional diagnostics.
Multi-threshold analysis (20, 30, 40, 50 cycles) confirms robustness: F1 scores remain above 88% across all thresholds, with the 30-cycle threshold achieving optimal 92.3% F1. At more conservative 20-cycle threshold, FNR drops to 0% (perfect recall) with acceptable 4.8% FPR increase, offering an alternative for ultra-conservative maintenance policies prioritizing absolute safety over cost efficiency.
V-E Performance Stratification by RUL Range
TABLE IV: RMSE by RUL Range (FD001 Best Run)
@@ V-E Performance Stratification by RUL Range
Table IV reveals performance stratification across degradation phases. As expected, prediction accuracy decreases in early operational life where degradation signatures are subtle and sensor readings remain near nominal values. However, as engines approach failure (RUL $\leq$ 30), our framework achieves exceptional accuracy (RMSE 5.14), precisely where accuracy matters most for preventing catastrophic failures while avoiding unnecessary premature maintenance.
@@
V-F Multi-Condition Dataset Performance
@@
V-G Ablation Study and Component Analysis
@@
V-H Computational Efficiency Analysis
@@ V-I Comparative Discussion
Computational Complexity vs. Deployment Feasibility: Transformer-based methods achieving 11-12 RMSE require 2M+ parameters and substantial computational resources. Domain-adaptation approaches like TMSCNN, while achieving excellent results, employ complex transfer learning pipelines. The 487K-parameter architecture achieves near 10ms inference on M1 hardware, making it practical for real-time onboard prognostics without ground station dependencies.
@@
V-J Practical Implications for Maintenance Scheduling